Becker CPA Review 2-7

Calculate the taxpayer's 2019 qualifying business income deduction for a qualified trade or business:

Filing status: Single
Taxable income: $180,000
Net capital gains: $0
Qualified business income (QBI): $80,000
W–2 wages: $20,000

a.$16,000

b.$10,000

c.$2,700

d.$13,684
Answer #1

Qualified business income deduction is the amount deducted from taxpayer's taxable income. It is calculated 20% of taxpayer's qualified business income.

Step 1.

20% of Qualified business income deduction ($80,000 × 20%)= $16,000

calculate the Wage-2 limit:

Greater of

50% of employee’s wages= ($20,000 × 50%)= $10,000

25% wages + 2.5% unadjusted basis immediately after acquisition= ($20,000×25% + 2.5%×0 )

=$ 5000

Step 2

If Wage-2 limit is less than the 20% of Qualified business income, the excess amount will be calculated:

20% of Qualified business income – Wage-2 limit= $16,000 - $10,000

                                                       Excess amount = $6,000

Step 3

Reduction Ratio for Single-Taxable income-$160700/$50000

=(180000-160700)/50000

=38.6%

Step 4- Qualified business income deduction allowed

=(0.2*QBI)-(Reduction Ratio * excess amount)

=(0.2*80000)-(38.6%*6000)

=16000-2316

=13684

Correct Answer is D

Please provide rating